146 g of water at 76.5 degree C are mixed with 75.
Answer #1

Q = m Cp dT formula

for first one :

m1 = 146 g

Cp = 4.184 J / oC g

dT = 76.5 - Tf

Q = m Cp dT = 146 x 4.184 (76.5 - Tf)

Q = 610.9 (76.5 - Tf ) --------------------------> heat loss

heat gain by the other water:

Q =75.4 x 4.184 x (Tf - 21.5)

Q = 315.5 (Tf - 21.5) --------------------------> heat gain

610.9 (76.5 - Tf )   = 315.5 (Tf - 21.5)

76.5 - Tf = 0.5164 (Tf -21.5)

76.5 - Tf = 0.5164 Tf - 11.1

87.6 = 1.5164 Tf

Tf = 57.8 oC

finla temperature = 57.8 oC